# DSwinIR: Rethinking Window-based Attention for Image Restoration

### v2

@@ IV-D Discussion and Limitation
The extensive experiments in this paper provide strong validation for our core thesis: a first-principles evolution of the backbone architecture is a direct and potent path to advancing all-in-one image restoration. Our proposed DSwinIR, built upon the novel Deformable Sliding Window attention, consistently established a new state-of-the-art across a wide range of benchmarks, often outperforming even sophisticated prompt-guided systems. This superiority is most pronounced in its massive performance leaps on tasks with spatially-variant degradations, such as dehazing and real-world deraining. We attribute this success to the synergistic combination of the token-centric sliding mechanism, which eradicates boundary artifacts, and the content-aware deformable sampling, which creates adaptive, non-rigid receptive fields. This unique capability allows DSwinIR to effectively model complex degradation patterns that are intractable for methods relying on static attention schemes, confirming the power of our architectural redesign.
Despite its demonstrated strengths, our work also illuminates several promising avenues for future research. First, while our architecture excels at restoring spatial structure, its performance on tasks dominated by extreme photometric adjustments, such as low-light enhancement, is competitive rather than leading. Future work could explore hybrid approaches that combine our powerful backbone with specialized modules for global color mapping. Second, our focus on architectural innovation means adapting DSwinIR to new tasks requires full retraining. A compelling direction is to integrate our backbone with parameter-efficient fine-tuning (PEFT) techniques, creating systems that are both powerful and highly adaptable. Finally, we believe the core principles of token-centric, content-adaptive attention hold significant potential beyond restoration, offering a promising new building block for other dense prediction tasks like semantic segmentation and object detection, where the ability to dynamically model features is paramount.
@@ V Conclusion
In this paper, we challenged the prevailing paradigm of window-based self-attention, arguing that its rigid, non-overlapping partitioning scheme imposes fundamental limitations on performance in image restoration. We introduced the Deformable Sliding Window Transformer (DSwinIR), a new foundational architecture built upon our novel DSwin attention mechanism. By pioneering a token-centric, content-aware paradigm that synergistically combines a sliding window to eliminate boundary artifacts with a deformable mechanism to learn adaptive receptive fields, DSwinIR achieves a more powerful and principled approach to visual feature representation. Extensive experiments rigorously demonstrated that this architectural evolution translates directly into substantial performance gains, establishing new state-of-the-art results across a diverse array of all-in-one and single-task restoration benchmarks. Ultimately, this work validates the critical importance of advancing fundamental architectures and offers a powerful new building block that we hope will inspire future developments in image restoration and beyond.

### v3

@@ IV-D Discussion and Limitation
The extensive experiment results support the view that redesigning window attention with sliding and deformable sampling is effective across various benchmarks. Despite its demonstrated strengths, our work also illuminates several promising avenues for future research. First, while our architecture excels at restoring spatial structure, its performance on tasks dominated by extreme photometric adjustments, such as low-light enhancement, is competitive rather than leading. A promising future direction is to integrate explicit global photometric modeling, for instance, by combining our powerful backbone with lightweight techniques like 3D Look-Up Tables (3D-LUTs) or by incorporating a learnable codebook prior. Second, our focus on architectural innovation means adapting DSwinIR to new tasks requires full retraining. A compelling direction is to integrate our backbone with parameter-efficient fine-tuning (PEFT) techniques, creating systems that are both powerful and highly adaptable. Finally, we believe the core principles of token-centric, content-adaptive attention hold significant potential beyond restoration, offering a promising new building block for other dense prediction tasks like semantic segmentation and object detection, where the ability to dynamically model features is paramount.
@@ V Conclusion
In this paper, we revisit window-based self-attention for image restoration and propose the Deformable Sliding Window (DSwin) attention and the DSwinIR backbone. The token-centric sliding window reduces boundary effects by centering each token within its local context, while deformable sampling adapts the receptive field to image content. Extensive evaluations on all-in-one and single-task benchmarks show consistent improvements, including state-of-the-art results on several datasets. We hope DSwin will serve as a useful building block for future image restoration architectures and related dense prediction tasks.
[Uncaptioned image] Gang Wu received the B.E. degree in the School of Computer Science and Technology from Soochow University, Jiangsu, China, in 2020. He is currently pursuing the Ph.D. degree in Faculty of Computing at Harbin Institute of Technology. His research interests include image restoration, representation learning, and self-supervised learning.
[Uncaptioned image] Junjun Jiang received the B.S. degree in Mathematics from the Huaqiao University, Quanzhou, China, in 2009, and the Ph.D. degree in Computer Science from the Wuhan University, Wuhan, China, in 2014. From 2015 to 2018, he was an Associate Professor with the School of Computer Science, China University of Geosciences, Wuhan. From 2016 to 2018, he was a Project Researcher with the National Institute of Informatics (NII), Tokyo, Japan. He is currently a Professor with the School of Computer Science and Technology, Harbin Institute of Technology, Harbin, China. His research interests include image processing and computer vision.
[Uncaptioned image] Kui Jiang received the M.E. and Ph.D. degrees from the School of Computer Science, Wuhan University, Wuhan, China, in 2019 and 2022, respectively. Before July 2023, he was a Research Scientist with the Cloud BU, Huawei. He is currently an Associate Professor with the School of Computer Science and Technology, Harbin Institute of Technology. He received the 2022 ACM Wuhan Doctoral Dissertation Award. His research interests include image/video processing and computer vision.
[Uncaptioned image] Xianming Liu received the B.S., M.S., and Ph.D. degrees in computer science from the Harbin Institute of Technology (HIT), Harbin, China, in 2006, 2008, and 2012, respectively. In 2011, he spent half a year at the Department of Electrical and Computer Engineering, McMaster University, Canada, as a Visiting Student, where he was a Post-Doctoral Fellow from 2012 to 2013. He was a Project Researcher with the National Institute of Informatics (NII), Tokyo, Japan, from 2014 to 2017. He is currently a Professor with the School of Computer Science and Technology, HIT. His research interests include trustworthy AI, 3D signal processing and biomedical signal processing.
[Uncaptioned image] Liqiang Nie received the B.Eng. and Ph.D. degree from Xi’an Jiaotong University and National University of Singapore (NUS), respectively. After Ph.D., he continued his research in NUS as a research fellow for three years. He is currently the dean of the Department of Computer Science and Technology, Harbin Institute of Technology (Shenzhen). His research interests lie primarily in multimedia computing and information retrieval. He is an AE of IEEE TIP, IEEE TKDE, IEEE TMM, IEEE TCSVT, ACM ToMM, and Information Science. Meanwhile, he serves as the chair of ICMR 2025, ICME 2025, and ACM MM 2027. He is a member of ICME steering committee. He has received many awards over the past four years, like SIGMM rising star in 2020, MIT TR35 China 2020, SIGIR best student paper in 2021, IEEE AI’s 10 to Watch in 2022, ACM MM Best paper award in 2022, and the National Youth Science and Technology Award in 2024.